OpenStudy (anonymous):

Use implicit differentiation to find dy/dx when 4y^2 = 3x-5/3x+5 I've done it but I keep getting the wrong answer, I must be doing something wrong but I don't know where

OpenStudy (anonymous):

\[4y ^{2}=\frac{ 3x-5 }{ 3x+5 }\]

jimthompson5910 (jim_thompson5910):

what are you getting so far?

OpenStudy (anonymous):

well if i'd type out all my working it'd be a bit tedious but my final answer i got was (-12y^2+3) / (24xy+40y)

jimthompson5910 (jim_thompson5910):

The first step is to derive both sides with respect to x. What do you get when you derive 4y^2 with respect to x?

OpenStudy (anonymous):

well first i brought the (3x+5) to the LHS and then actually expanded it, am I allowed to expand it? so I derived 12xy^2 actually not 4y^2

jimthompson5910 (jim_thompson5910):

If you did that, you would get 4y^2(3x+5) = 12xy^2+20y^2 Then you would derive 12xy^2+20y^2

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

giving you what?

OpenStudy (anonymous):

12y^2 + 24xy dy/dx + 40y dy/dx

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now derive the RHS 3x-5

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so then so far it would look like this, well this is what i got. 12y^2 + 24xy dy/dx + 40y dy/dx = 3

jimthompson5910 (jim_thompson5910):

the last thing to do is isolate dy/dx

OpenStudy (anonymous):

by taking it as a common factor from 24 and 40 terms

OpenStudy (anonymous):

dy/dx (24xy + 40y) = -12y^2 + 3 dy/dx = -12y^2 + 3 / 24xy + 40y

OpenStudy (anonymous):

but that's not the correct answer so I'm not sure where I went wrong?

jimthompson5910 (jim_thompson5910):

I think you meant to say \[\Large \frac{dy}{dx} = \frac{-12y^2+3}{24xy+40y}\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Keep in mind that initially \[\large 4y^{2}=\frac{ 3x-5 }{ 3x+5 }\] so how can we use that -12y^2 to replace it with something in terms of x?

OpenStudy (anonymous):

sorry i'm not too sure...

jimthompson5910 (jim_thompson5910):

notice how -12y^2 = -3*4y^2

OpenStudy (anonymous):

yeh

jimthompson5910 (jim_thompson5910):

so you can replace 4y^2 with the RHS of \[\large 4y^{2}=\frac{ 3x-5 }{ 3x+5 }\]

jimthompson5910 (jim_thompson5910):

and simplify

OpenStudy (anonymous):

wait u mean, u can substitute 4y^2 = 3x-5 / 3x + 5 into our final dy/dx equation as 4y^2 right? then simplify?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

it's going to be a bit messy, but it can be done

OpenStudy (anonymous):

is this step correct? 30 / (3x+5) (24xy + 40y)

jimthompson5910 (jim_thompson5910):

if 24xy+40y is in the denominator, then yes it's correct

jimthompson5910 (jim_thompson5910):

you can simplify further

OpenStudy (anonymous):

yeh, I just hate going this much in depth just for simplifying but i think i'm getting there.

jimthompson5910 (jim_thompson5910):

tell me what you get

jimthompson5910 (jim_thompson5910):

what does 24xy + 40y factor to?

OpenStudy (anonymous):

8y(3x+5)

jimthompson5910 (jim_thompson5910):

So (3x+5)(24xy + 40y) turns into (3x+5)*8y(3x+5) = 8y(3x+5)^2

jimthompson5910 (jim_thompson5910):

Giving you \[\large \frac{ 30 }{ 8y(3x+5)^2 }\] \[\large \frac{ 15 }{ 4y(3x+5)^2 }\]

OpenStudy (anonymous):

yep k, i got it.

OpenStudy (anonymous):

thanks so much jim!

OpenStudy (anonymous):

so in the end I was doing the right steps, just didn't simplify doing further like 100 steps...

jimthompson5910 (jim_thompson5910):

here's a faster/simpler way to do it (in my opinion anyway) \[\large 4y^{2}=\frac{ 3x-5 }{ 3x+5 }\] \[\large \frac{d}{dx}[4y^{2}]=\frac{d}{dx}\left[\frac{ 3x-5 }{ 3x+5 }\right]\] \[\large 8y*\frac{dy}{dx}=\frac{d}{dx}\left[\frac{ 3x-5 }{ 3x+5 }\right]\] \[\large 8y*\frac{dy}{dx}=\frac{ 30 }{ (3x+5)^2 } \ ... \ \text{See note in next post}\] \[\large \frac{dy}{dx}=\frac{ 30 }{ 8y(3x+5)^2 }\] \[\large \frac{dy}{dx}=\frac{ 15 }{ 4y(3x+5)^2 }\]

jimthompson5910 (jim_thompson5910):

Deriving \[\Large \frac{ 3x-5 }{ 3x+5 }\] ---------------------------------------------------------------------------------- \[\Large h(x) = \frac{ f(x) }{ g(x) } = \frac{ 3x-5 }{ 3x+5 }\] \[\Large f(x) = 3x-5\] \[\Large g(x) = 3x+5\] \[\Large f^{\prime}(x) = 3\] \[\Large g^{\prime}(x) = 3\] ---------------------------------------------------------------------------------- Quotient Rule: \[\Large h(x) = \frac{ f(x) }{ g(x) }\] \[\Large h^{\prime}(x) = \frac{ f^{\prime}(x)g(x) - g^{\prime}(x)f(x) }{ [g(x)]^2 }\] \[\Large h^{\prime}(x) = \frac{ 3(3x+5)-3(3x-5) }{ (3x+5)^2 }\] \[\large h^{\prime}(x)=\frac{ 9x+15-9x+15 }{ (3x+5)^2 }\] \[\large h^{\prime}(x)=\frac{ 30 }{ (3x+5)^2 }\]

OpenStudy (anonymous):

yeh that is much faster! ==

OpenStudy (anonymous):

well Jim, thanks so much for ur help, guidance, tips and tricks! hopefully u'll be able to help me more in the future when u have time lol

jimthompson5910 (jim_thompson5910):

you're welcome